OpenStudy (anonymous):

Solve the system of equations. Show work Please -2x – 1 = y x = -y + 3

OpenStudy (anonymous):

multiply bottom times 2, then add both equations together. you will be left with only y, isolate y. then repeat process from the beginning for x

OpenStudy (anonymous):

\[\large -2x – 1 = y[1]\] \[\large x = -y + 3 [2]\] The first equation tells you that "y" can be written as -2x-1. So you rewrite y in the second equation as -2x-1. \[\large x=-(-2x-1)+3\] That's your y in the brackets. Expand it out or in other words, distribute. \[\large x=--2x--1+3\] Two minus signs make a plus sign. \[\large x=+2x+1+3\] Don't need the plus sing in the very front because it doesn't look good. \[\large x=2x+1+3\] Gather like terms. \[\large x=2x+4\] \[\large x+4=0\] Move the 4 to the other side. Remember to change the sign of a term when it moves to the other side of the equal sign. \[\large x=-4\] TO find y, just easily put the x-value into the first equation. \[\large y=-2x-1\] \[\large y=-2(-4)-1\] \[\large y=8-1\] \[\large y=7\]